# ReadMe.LLM: A Framework to Help LLMs Understand Your Library

### v2

@@ Abstract
Large Language Models (LLMs) often struggle with code generation tasks involving niche software libraries. Existing code generation techniques with only human-oriented documentation can fail —– even when the LLM has access to web search and the library is documented online. To address this challenge, we propose ReadMe.LLM, LLM-oriented documentation for software libraries. By attaching the contents of ReadMe.LLM to a query, performance consistently improves to near-perfect accuracy, with one case study demonstrating up to 100% success across all tested models. We propose a software development lifecycle where LLM-specific documentation is maintained alongside traditional software updates. In this study, we present two practical applications of the ReadMe.LLM idea with diverse software libraries, highlighting that our proposed approach could generalize across programming domains.
@@ 1 Introduction
Large Language Models (LLMs) like GPT [1] and Llama [2] have transformed the software development ecosystem. More engineers are using LLMs to generate code with existing software libraries, leveraging these tools to approach coding tasks more efficiently and intuitively. In some cases, we are even seeing AI agents begin to replace human developers themselves. These models, often used as coding assistants, are capable of generating code, debugging, and creating documentation through natural language prompting. One emerging trend, frequently referred to as “vibe coding” [3], involves engineers prompting LLMs with simple, high-level natural language instructions and iteratively refining their code based on the model’s suggestions. This interactive exploratory approach enables fast prototyping and creates a more fluid software development process.
@@ 1.2 Existing Solutions
There are different techniques for leveraging LLMs for code generation tasks. We list key categories and examples in Figure 1. Many strategies, such as Retrieval Augmented Generation (RAG) [9], require additional infrastructure that falls outside of the typical developer workflow using standard IDEs111Of course, one possible view of the future would involve building a well standardized approach to integrating RAG with IDEs and copilots [10]. But we are not there yet.. The most accessible approach that offers the least friction to development is Prompt Engineering [11]. Within Prompt Engineering, there has been research on thought generation, such as Chain of Thought [12], or how models perform in a “Zero-Shot” [13] manner. We propose a specific prompting framework -– ReadMe.LLM —-which leverages assets (e.g. function signatures, examples, and descriptions) from a respective software library to assist code generation tasks.
@@
Current documentation, such as ReadMe.md files, is written for human readers, but LLMs interpret information differently and are less effective with human-targeted formats. We argue that there should be LLM-targeted documentation. To address this, we propose ReadMe.LLM, a structured format to streamline library usage by LLMs:
@@
This was the structure we found worked best based on the libraries we experimented with; however, libraries from different domains may need to make small adjustments. We used XML tags to separate different types of content (e.g. examples). This formatting improves readability for LLMs and helps them easily parse the rules, description, and code snippets in ReadMe.LLM [23].
@@ 3.3 Held Out Tests
To further assess the robustness of ReadMe.LLM, we conducted a held-out test using two new models that had not been used in previous experiments: Gemini 2.0 Flash and Mistral Large. This evaluation aimed to verify the effectiveness of our final framework when applied to LLMs that did not contribute to our development process.
We began with Supervision. Even with web search enabled, zero-context prompting performed poorly. Gemini succeeded only once out of five trials on the first task, and both models failed entirely on the second. With ReadMe.LLM Gemini’s success rate jumped to 100% on the first task and 80% on the second. Mistral, which had previously failed both, reached a perfect 100% on both tasks. Without web search, Supervision’s ReadMe.LLM sustained a high success rate with both models, with only Gemini exhibiting a slight decline.
We tested these same models with DigitalRF. When prompted without any additional context, accuracy was 0%, even with web search capabilities. Once ReadMe.LLM was applied, however, both models achieved an 80% success rate, showing a dramatic and consistent improvement. In line with the original models, the absence of web search capabilities resulted in a slight performance decline, but still remained significantly superior to conditions without context.
@@ 4 Discussion
It’s Possible to Seamlessly Improve Code Generation Through Prompting with LLM-Oriented Documentation. Through our experimentation and research, we show that prompting with LLM-oriented software library documentation —ReadMe.LLM– can greatly increase the performance of code completion tasks for LLMs. ReadMe.LLM can perform optimally in the majority of scenarios irrespective of model selection. This is a powerful tool for library developers and engineers, because it increases the accessibility and performance of leveraging LLMs for code completion. Engineers no longer have to create complex prompts, use computationally-intensive methods like RAG, or drastically change their queries. Instead, they can just attach a library’s ReadMe.LLM and ask questions as usual. Similarly, library developers can develop a ReadMe.LLM to ensure that their library is being correctly represented by an LLM and therefore, seamlessly used by the targeted engineer.
@@
Second, the model would often hallucinate by using an alternative method or library, failing to use the targeted library’s code. For example, with Supervision, the model would often crop images by slicing a dictionary, rather than using the crop image function defined in Supervision. This highlights the importance of adding examples and function definitions, especially when the functionality may conflict with existing libraries that a model is trained on. This can help ground the model in using the correct function from the targeted library.
@@ 5 Conclusion and Future Work
We present ReadMe.LLM, novel LLM-oriented documentation that provides relevant context about a software library to assist code generation. We evaluated different combinations and structures of context and tested these across the current leading LLMs. We presented the optimal ReadMe.LLM structure, which has the highest average accuracy across different models, and increases correctness by 5x.

### v3

@@ Abstract
Large Language Models (LLMs) often struggle with code generation tasks involving niche software libraries. Existing code generation techniques with only human-oriented documentation can fail – even when the LLM has access to web search and the library is documented online. To address this challenge, we propose ReadMe.LLM, LLM-oriented documentation for software libraries. By attaching the contents of ReadMe.LLM to a query, performance consistently improves to near-perfect accuracy, with one case study demonstrating up to 100% success across all tested models. We propose a software development lifecycle where LLM-specific documentation is maintained alongside traditional software updates. In this study, we present two practical applications of the ReadMe.LLM idea with diverse software libraries, highlighting that our proposed approach could generalize across programming domains.
@@ 1 Introduction
Large Language Models (LLMs) like GPT [1] and Llama [2] have transformed the software development ecosystem. More engineers are using LLMs to generate code with existing software libraries, leveraging these tools to approach coding tasks more efficiently and intuitively. In some cases, we are even seeing AI agents begin to replace human developers themselves. These models, often used as coding assistants, are capable of generating code, debugging, and creating documentation through natural language prompting.
These advances build on a growing lineage of code-specific language models. OpenAI’s Codex [3], fine-tuned on GitHub code, enabled natural language-to-code translation and powers tools like GitHub Copilot [4]. Meta’s Code LLaMA [5] was pretrained specifically on code-related data, allowing it to support multiple programming languages and longer context windows. Both models expanded the capabilities of LLMs while maintaining low-friction interfaces for developers — but they also highlight a core challenge: for this innovation and performance to continue, LLMs rely on consistent and structured documentation. Recent work has emphasized that structured specification (precise descriptions of a component’s expected behavior, inputs, and outputs) is essential to making LLM-based systems more modular and reliable [6]. As these systems become increasingly integrated into IDEs and developer workflows, new coding practices continue to emerge.
One emerging trend, frequently referred to as “vibe coding” [7], involves engineers prompting LLMs with simple, high-level natural language instructions and iteratively refining their code based on the model’s suggestions. This interactive exploratory approach enables fast prototyping and creates a more fluid software development process.
@@ 1.2 Existing Solutions
There are different techniques for leveraging LLMs for code generation tasks. We list key categories and examples in Figure 1. Many strategies, such as Retrieval Augmented Generation (RAG) [13], require additional infrastructure that falls outside of the typical developer workflow using standard IDEs111Of course, one possible view of the future would involve building a well standardized approach to integrating RAG with IDEs and copilots [14]. But we are not there yet.. The most accessible approach that offers the least friction to development is Prompt Engineering [15]. Within Prompt Engineering, there has been research on thought generation, such as Chain of Thought [16], or how models perform in a “Zero-Shot” [17] manner. CodeScholar [24] addresses this by generating realistic, idiomatic API usage examples to help developers understand how to use unfamiliar libraries, serving as a complement to traditional documentation. These examples not only improve human comprehension but also improve the quality of LLM-generated code when used in downstream tasks. We propose a specific prompting framework -– ReadMe.LLM —-which leverages assets (e.g. function signatures, examples, and descriptions) from a respective software library to assist code generation tasks.
@@
Current documentation, such as ReadMe.md files, is written for human readers, but LLMs interpret information differently and are less effective with human-targeted formats. We argue that there should be LLM-targeted documentation 222As we were writing this report after completing our experiments, Andrej Karpathy tweeted: ”It’s 2025 and most content is still written for humans instead of LLMs. 99.9% of attention is about to be LLM attention, not human attention. E.g. 99% of libraries still have docs that basically render to some pretty .html static pages assuming a human will click through them. In 2025 the docs should be a single your_project.md text file that is intended to go into the context window of an LLM. Repeat for everything.”[28].. To address this, we propose ReadMe.LLM, a structured format to streamline library usage by LLMs:
@@
This was the structure we found worked best based on the libraries we experimented with; however, libraries from different domains may need to make small adjustments. We used XML tags to separate different types of content (e.g. examples). This formatting improves readability for LLMs and helps them easily parse the rules, description, and code snippets in ReadMe.LLM [29]. The complete ReadMe.LLM docs used in our experiments can be found in Appendix C. In order to create a ReadMe.LLM for a library follow the instructions in Appendix B.
@@ Results
A comparison of results for earlier model versions across all contexts is provided in Appendix A
@@ 3.3 Held Out Tests
To further assess the robustness of ReadMe.LLM, we conducted a held-out test using three new models that had not been used in previous experiments: Gemini 2.0 Flash, Mistral Large, and GPT-4.1. This evaluation aimed to verify the effectiveness of our final framework when applied to LLMs that did not contribute to our development process.
We began with Supervision. Even with web search enabled, zero-context prompting performed poorly. With ReadMe.LLM Gemini’s success rate jumped to 100% on the first task and 80% on the second. Mistral, which had previously failed both, reached a perfect 100% on both tasks, as did GPT-4.1. Without web search, Supervision’s ReadMe.LLM sustained a high success rate with these models, with only Gemini exhibiting a slight decline.
We tested these same models with DigitalRF. When prompted without any additional context, accuracy was 0%, even with web search capabilities. However, once ReadMe.LLM was applied, Gemini and Mistral achieved a 80% success rate and GPT-4.1 100%, showing a dramatic and consistent improvement. In line with the original models, the absence of web search capabilities resulted in a slight performance decline with Gemini and Mistral, but still remained significantly superior to conditions without context.
@@ 4 Discussion
It’s Possible to Seamlessly Improve Code Generation Through Prompting with LLM-Oriented Documentation. Through our experimentation and research, we show that prompting with LLM-oriented software library documentation —ReadMe.LLM– can greatly increase the performance of code completion tasks for LLMs. ReadMe.LLM can perform optimally in the majority of scenarios irrespective of model selection. This is a powerful tool for library developers and engineers, because it increases the accessibility and performance of leveraging LLMs for code completion. Engineers no longer have to create complex prompts, use computationally-intensive methods like RAG, or drastically change their queries. Instead, they can just copy and paste a library’s ReadMe.LLM into an LLM’s prompt window and give a code generation task as usual 333After our paper was first released, Andrej Karpathy tweeted in support of the seamless experience of copy and pasting documentation for LLMs: ”Tired: elaborate docs pages for your product/service/library with fancy color palettes, branding, animations, transitions, dark mode, … Wired: one single docs .md file and a ’copy to clipboard’ button” [47].. Similarly, library developers can develop a ReadMe.LLM to ensure that their library is being correctly represented by an LLM and therefore, seamlessly used by the targeted engineer.
@@
Second, some models are able to complete the task but do not use the functions provided by the specified library. We marked these occurrences as failures, since utilizing the actual library functions is typically much more efficient than assembling a solution from a mix of other libraries. For example, during our Supervision tasks, the library offered functions to crop or overlay an image, but the LLM generated code that performed these tasks using the CV library instead, resulting in much longer and less efficient code. This is also important from a Software Bill of Materials (SBOM) [48] perspective: an SBOM is an inventory of all software components and dependencies, and using the correct library functions ensures that the SBOM accurately reflects the software’s dependencies, improving transparency and security. This highlights the importance of providing clear examples and function definitions, especially when the desired functionality may overlap with existing libraries that a model is trained on. Doing so can help guide the model to use the correct function from the targeted library.
LLM Performance is a Moving Target. While our final results demonstrate consistent gains using ReadMe.LLM, earlier versions of the same models – shown in Appendix A – showed far more dramatic improvements.
This raises a broader question: as models evolve, which challenges are temporary and which are more fundamental? For example, hallucinations might fade as models improve, or they might persist in new forms. Recent work from Nexusflow on Athene-V2 suggests that models often face a tradeoff: being fine-tuned for conversational chat vs. agentic behavior [49]. This specialization means models may not be able to excel at all tasks simultaneously. And regardless of progress, smaller local models will likely continue to face more limitations due to resource constraints. All of this reinforces the idea that LLM evaluation must keep pace with a fast-moving target.
Jailbreaking and Prompt Robustness. Recent safety research underscores that LLM behavior can be dramatically altered through structured prompting alone. For instance, HiddenLayer’s KROP technique bypasses safety safeguards simply by reformatting the input using specific structural patterns, without needing to inject unsafe content [50]. Inspired by this behavior, ReadMe.LLM encloses library information in XML tags – a way that LLMs are highly responsive to. However, this strength may also be a weakness: such formatting has recently been characterized as a form of ”jailbreaking,” since many models appear overly obedient to syntactically clean inputs like XML. If future alignment interventions are trained to block these structured jailbreaks, ReadMe.LLM’s effectiveness could degrade – even if its content remains entirely safe and constructive. This highlights the need to monitor prompt performance over time and investigate how safety tuning may unintentionally interfere with developer-facing tools.
Similar Approaches in Practical Evaluation. The Berkeley Function-Calling Leaderboard has implemented data filtering and quality enhancement techniques when using live, user-contributed function documentation and queries to benchmark LLM performance in function calling [51, 52]. Their definition of high-quality documentation—structured JSON containing key fields such as function name, description, and detailed specifications—closely aligns with the design of ReadMe.LLM, proving the generalizability of our structure. Furthermore, their evaluation includes comprehensive benchmarking across programming languages and libraries, incorporating methods such as function relevance detection and Abstract Syntax Tree (AST) analysis. Integrating these evaluation techniques into the ReadMe.LLM assessment could enhance its robustness to benchmark and highlight its practical benefits.
In effect, the new thing we are advocating is that every library developer should learn from the benchmark design, just as LLM developers learn from benchmarks. Knowing that LLM developers are chasing higher benchmark scores also informs library developers. While perhaps large popular libraries have the luxury of expecting LLM-providers to optimize for them, smaller niche libraries would do well to make themselves look like the benchmarks that LLM providers optimize for.
Extending ReadMe.LLM to Tool Use. Recent work has increasingly focused on LLMs’ ability to interact with external tools, such as invoking APIs or executing functions. As LLM applications become more powerful and autonomous, their ability to correctly use APIs becomes critical. Similar to software libraries, APIs define the correct way to access desired functionality through structured calls. Inspired by this similarity, we propose extending ReadMe.LLM to support tool use by generating LLM-targeted documentation for APIs. Specifying important details like endpoints, parameters, or expected responses can be formatted and optimized for LLMs. Notably, the API-Bank benchmark demonstrates the importance of LLMs understanding API structure and semantics to improve tool use accuracy [53]. ReadMe.LLM can bridge the current performance gaps by providing LLM-friendly API documentation.
@@ 5 Conclusion and Future Work
We present ReadMe.LLM, novel LLM-oriented documentation that provides relevant context about a software library to assist code generation. We evaluated different combinations and structures of context and tested these across the current leading LLMs. We presented the optimal ReadMe.LLM structure, which has the highest average accuracy across different models, and increases correctness up to 100%.
@@ Acknowledgments
Appendix A Appendix: Comparison of results with older models
A.1 DigitalRF’s Case Study with Older Models
In our original experiments on the DigitalRF case study, we evaluated Claude 3.5 Sonnet across all eight context combinations and ReadMe.LLM, with web search enabled. When Claude 3.7 Sonnet was released, we repeated the same experiments.
A clear pattern emerges: while both versions benefit from richer context, the older Claude 3.5 model exhibited a larger performance jump when provided with our ReadMe.LLM versus human-oriented documentation. Specifically, Claude 3.5’s accuracy moved from 20% (no context or ReadMe.md) to 80% with ReadMe.LLM—a 60% point gain—whereas Claude 3.7 improved from 60% (no context or ReadMe.md) to 100% with ReadMe.LLM—a 40% point gain. This indicates that ReadMe.LLM had a greater impact in bridging the knowledge gap in the earlier version of the model. As models become more capable (e.g., Claude 3.7), the baseline performance on human-oriented docs rises, but ReadMe.LLM continues to provide a consistent boost.
When a new Sonar model was released built on Llama 3.3 70B, we also repeated our DigitalRF experiments to compare it with the older Sonar based on Llama 3.1 70B. Although we did not evaluate ReadMe.LLM on the older Sonar (Llama 3.1 70B), the eight standard context combinations reveal some clear trends. Both Sonar 3.1 and Sonar 3.3 achieved 100% success when provided only with examples, underscoring the power of real-world snippets. Most notably, Sonar 3.1 hallucinated dramatically under richer prompts—success falls to 40% with ReadMe.md + Examples and 20% with the full mixed context—whereas Sonar 3.3 remained far more robust (100% and 80%, respectively). In both versions, ReadMe.md alone yielded only a 20% success rate, again showing that human-oriented documentation offers minimal benefit.
A.2 Supervision’s Case Study with Older Models
When Claude was upgraded from 3.5 Sonnet to 3.7 Sonnet, we also reran Supervision’s Task 1 case study experiments. Claude 3.5 failed always with no context or ReadMe.md alone, yet jumped to a perfect 100% as soon as it saw examples, functions, or any combined context—including ReadMe.LLM. In contrast, Claude 3.7 already achieved 60% with no context and 20% with ReadMe.md, and likewise hits 100% once provided with ReadMe.LLM. Similar to the point we made for DigitalRF’s case study, the improvement from human-oriented docs to ReadMe.LLM is far more dramatic for Claude 3.5 than for Claude 3.7.
Appendix B Appendix: ReadMe.LLM Instructions
1.
Determine Functions to Include: Select the most popular or important functions and classes that represent the library’s core purpose and domain context.
2.
Extract Function Signatures: Clearly include function and method definitions with parameters and return types
3.
Find/Create Code Examples of Each Function: Pair signatures with illustrative code examples that demonstrate real-world usage of the functions
4.
Interleave Description of Function, Function Signature, and Example: Structure the content using sections with XML tags, combining a description, the function signatures, and the examples in a parsable layout. Below is the structure of the core building block of our ReadMe.LLM:
<context_1> <context_1_description> </context_1_description> <context_1_function> </context_1_function> <context_1_example> </context_1_example> </context_1>
5.
Add rules to the top of the ReadMe.LLM file. Below are the rules we added to our ReadMe.LLM files:
Rules:
Rule number 1: When you’re unsure about something, ask the user what information you need.
Rule number 2: Reuse SuperVision’s functions and code when applicable.
Rule number 3: Consider library dependencies when generating code solutions.
Appendix C Appendix: ReadMe.LLM Content
C.1 Supervision’s ReadMe.LLM
The following is the full ReadMe.LLM for the Supervision library:
C.2 DigitalRF’s ReadMe.LLM
The following is the full ReadMe.LLM for the DigitalRF library: